Over the last four weeks, the stock price for Royal Caribbean has closed at $40.00, $44.00, $40.00, and $36.00, respectively. Based on these prices, what is the expected return and standard deviation of this stock? (Assume that Royal Caribbean has not paid any dividends during this period.)
Answer #1

Expected return=(40/44-1+44/40-1+40/36-1)/3=4.007%

Standard Deviation=sqrt((40/44-1-4.007%)^2+(44/40-1-4.007%)^2+(40/36-1-4.007%)^2)/sqrt(3-1)=11.356%